OpenStudy (anonymous):

Can anyone help with Geometry Semester 1 in apex virtual school?

OpenStudy (anonymous):

Whats the question?

OpenStudy (anonymous):

its a quiz :/

OpenStudy (anonymous):

tan(a)=7/x and also tan(a)=84/(156-x) can you solve it now?

OpenStudy (anonymous):

im not good with equations!

OpenStudy (anonymous):

someone help :X

OpenStudy (anonymous):

Is this freshman geometry?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

7/x=84/(156-x) 7(156-x)=84x 156-x=12x 13x=156 x=156/13 x=12

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

;)